If a triangle has a height of 12 in. and a base of 5 in., what is its area?
Answer by nyc_function(2741)

A = 1/2 times base times height
A = 1/2 times 12in times 5in
A = 1/2 times 12in times 5in
A = 1/2 times 60in
A = 30in^2